OpenStudy (anonymous):

(4x-3y) ^3 expand the binomial. help :D

OpenStudy (aravindg):

Use binomial expansion!

OpenStudy (anonymous):

4x^3-3y^0 + 4x^2-3y^1+4x^1-3y2?

OpenStudy (mathmale):

You have several choices here. You could memorize the first few binomial expansions: (a+b)=a+b (a+b)^2 = a^2 + 2ab + b^2 (a+b)^3= a^3 + 3a^2b +3ab^2 + b^3 or you could multiply out (a+b)(a+b)(a+b), or you could look up and apply Pascal's Triangle. Take your pick.

OpenStudy (mathmale):

Please note: 4x^3 is not the same as the required (4x)^3.

OpenStudy (anonymous):

Death! I am ripping my eyes out. Melted brains drip out of empty sockets onto a bed of burning coals.

OpenStudy (mathmale):

You've just earned an "A" for drama. But seriously, how could I best help you?

OpenStudy (anonymous):

You have : ) I think I just need some sleep. I'll come back to this tomorrow morning.